# An EPTAS for multiprocessor scheduling with rejection under a machine cost constraint

### v1

@@ 4.2 Restricted instance
There are at most $(1+\frac{\epsilon}{2})\frac{m}{\Delta}$ large jobs, i.e., $|L|\leq(1+\frac{\epsilon}{2})\frac{m}{\Delta}$.
@@
By Eq.(9), we have $(1+\frac{\epsilon}{2})\frac{1}{\Delta}=\frac{1}{\Delta}+2m+6$, which is an integer. Assume there are at least $(1+\frac{\epsilon}{2})\frac{m}{\Delta}+1$ large jobs and thus in the schedule $\pi^{\prime}$, there exists a machine where at least $(1+\frac{\epsilon}{2})\frac{1}{\Delta}+1$ large jobs are scheduled in $\pi^{\prime}$. By Eq.(8), the makespan of $\pi^{\prime}$ is at least $((1+\frac{\epsilon}{2})\frac{1}{\Delta}+1)\Delta>1+\frac{\epsilon}{2}\geq(1+\frac{\epsilon}{2})Z^{*}(I)$, which contradicts Lemma 15. ∎
Note that each large job must be assigned to some machine in $\pi^{\prime}$. The following definition describes all possible assignments for large jobs.
Definition 18.
(Large job assignment types) A large job assignment type is an $|L|$-tuple $(\ell_{1},\ldots,\ell_{|L|})$ such that each $\ell_{i}\in\{1,2,\ldots,m\}$, for every $i=1,\ldots,|L|$.
A schedule is associated with a large job assignment type $(\ell_{1},\ldots,\ell_{|L|})$ if the $i$-th large job is assigned to the machine $M_{\ell_{i}}$ for every $i=1,\ldots,|L|$.
@@
The number of distinct large job assignment types is at most $m^{(1+\frac{\epsilon}{2})\frac{m}{\Delta}}$, which is a fixed constant when $m$ is a constant.
@@
Similarly to Lemma 14, the lemma follows from $\ell_{i}\in\{1,2,\ldots,m\}$ and $|L|\leq(1+\frac{\epsilon}{2})\frac{m}{\Delta}$. ∎
Clearly, the schedule $\pi^{\prime}$ is associated with a unique large job assignment and by Lemma 19, we are able to enumerate all possible large job assignments to find the one associated with $\pi^{\prime}$. At this time, we have determined the rejection decision for each risky job and also determined the schedule of each large job.
4.3 Schedule tiny and safe jobs
We remind the readers that all tiny jobs are accepted by our algorithm. In this section, we present a linear program to determine the schedule of most tiny jobs and; determine the rejection decisions of most safe jobs and the designated machine for each accepted safe job. For every safe job $J_{j}\in S$, let $x_{j}\in\{0,1\}$ be a binary decision variable such that $x_{j}=1$ if and only if $J_{j}$ is rejected in the solution $\pi^{\prime}$. Moreover, for every (tiny or safe) job $J_{j}\in T\cup S$ and every machine $M_{i}$, let $y_{ji}\in\{0,1\}$ be a binary decision variable, which is $1$ if and only if $J_{j}$ is scheduled on the machine $M_{i}$ in $\pi^{\prime}$. Recall that we schedule all large jobs according to a given large job assignment type and we let $t_{i}$ be the total processing time of large jobs assigned to the machine $M_{i}$, for every $i=1,\ldots,m$. We are now ready to present our linear program.
$\min$ $C_{\max}+\sum_{J_{j}\in S}x_{j}e_{j}$ (11)
$\sum_{i=1}^{m}y_{ji}=1$ $\forall J_{j}\in T$
$x_{j}+\sum_{i=1}^{m}y_{ji}=1$ $\forall J_{j}\in S$
$C_{\max}\geq t_{i}+\sum_{J_{j}\in T\cup S}y_{ji}p^{\prime}_{j}$ $\forall M_{i}$
$\sum_{J_{j}\in S}(1-x_{j})p_{j}\leq U-\sum_{J_{j}\in A^{\prime}_{1}}p^{\prime}_{j}.$
We next interpret the LP(11) in details. The first constraint indicates that every tiny job should be scheduled on some machine. Note that if a safe job $J_{j}$ is rejected, then $x_{j}=1$ and $y_{ji}=0$ for every machine $M_{i}$. Otherwise, $J_{j}$ is accepted and should be scheduled on some machine. Therefore, $x_{j}=0$ and exactly one $y_{ji}$ equals to $1$. In both cases, the second constraint of LP(11) is always satisfied. The third constraint is from the definition of makespan. By Remark 10, we know $p^{\prime}_{j}=p_{j}$ for every job. The last constraint indicates the total processing time of accepted jobs is bounded by $U$ and the jobs in $A^{\prime}_{1}$ has been accepted. Note that LP(11) has exactly $(m+1)|S|+m|T|\leq(m+1)n$ variables and $|T|+|S|+m+1\leq n+m+1$ constraints. The linear program is polynomial-time solvable in $n,m$ because the time complexity of a linear program is polynomial in the number of continuous variables and polynomial in the number of constraints [13].
We consider a basic solution of LP(11) and partition the tiny jobs into two sets as follows.
$T_{1}=\{J_{j}\in T:\mbox{ $y_{ji}=1$ for some $i$}\}\mbox{ and }T_{2}=T\setminus T_{1}.$ (12)
Similarly, we partition the safe jobs into three sets as follows.
$S_{1}=\{J_{j}\in S:\mbox{ $y_{ji}=1$ for some $i$}\},S_{2}=\{J_{j}\in S:\mbox{$x_{j}=1$}\},\mbox{ and }S_{3}=S\setminus(S_{1}\cup S_{2}).$ (13)
Lemma 20.
$|T_{2}|+|S_{3}|\leq m+1$.
@@ 4.3 Schedule tiny and safe jobs
Since the number of constraints of LP(11) is $|T|+|S|+m+1$, there are at most $|T|+|S|+m+1$ positive values in the basic solution. By Eqs.(12) and (13), it is easy to see the number of positive values is at least $|T_{1}|+2|T_{2}|+|S_{1}|+|S_{2}|+2|S_{3}|=|T|+|S|+|T_{2}|+|S_{3}|$. Therefore, we have $|T|+|S|+|T_{2}|+|S_{3}|\leq|T|+|S|+m+1$ and thus the lemma is proved. ∎
We accept the jobs in $S_{1}$ and assign each job $J_{j}$ in $T_{1}\cup S_{1}$ to the machine $M_{i}$ such that $y_{ji}=1$. Finally, we reject all safe jobs in $S_{2}\cup S_{3}$ and for the tiny job set $T_{2}$, we arbitrarily append at most two of them at the end of the sub-schedule on each machine. By Lemma 20, $|T_{2}|\leq m+1$ and therefore, all jobs of $T_{2}$ have been assigned to the machines. We present the high-level description of the whole algorithm as follows.
Algorithm 2 Algorithm Approx2 for $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
Input: An instance $J=\{J_{1},\ldots,J_{n}\}$, a fixed number $m$ of machines and a fixed constant $10<\epsilon<$ 1;
Call Approx1 in Algorithm 1 on the instance $I$ to obtain cost $Z\leq 2Z^{*}$;
use $Z$ to normalize instance $I$ so that $\frac{1}{2}\leq Z^{*}\leq Z=1$ (Eq.(8));
set up $\Delta$ using $\epsilon$ and $m$ as in Eq.(9) and partition jobs into risky and safe as in Definition 7;
set up $\delta$ using $\epsilon$ and $\Delta$ as in Eq.(10) and construct the restricted instance $I^{\prime}$as in Definition 9;
partition the risky jobs in $I^{\prime}$ into at most $\frac{1}{\delta}$ subsets $B_{i}$ as in Definition 12;
for each rejection type $(k_{1},\ldots,k_{1/\delta})$ as in Definition 13 do
     reject the first $k_{i}$ jobs in $B_{i}$ for every $i=1,\ldots,1/\delta$ and accept the other risky jobs;
     partition the accepted risky jobs into large and tiny as in Definition 16;
     for each large job assignment as in Definition 18 do
         assign each large job to the machine stated by the large job assignment;
         construct the linear program as in LP(11) and compute a basic solution;
         partition the tiny jobs into two sets $T_{1},T_{2}$ as in Eq.(12)
         partition the safe jobs into three sets $S_{1},S_{2},S_{3}$ as in Eq.(13);
         accept the jobs in $S_{1}$ and assign each job $J_{j}$ in $T_{1}\cup S_{1}$ to the machine $M_{i}$ such that $y_{ji}=1$;
         reject the jobs in $S_{2}\cup S_{3}$;
         append at most two jobs in $T_{2}$ at the end of the sub-schedule on each machine;
     end for
end for
return the feasible solution with the minimum cost for $I^{\prime}$;
replace back the original rejection penalty of each job in the selected solution.
Theorem 21.
Algorithm Approx2 is an EPTAS for $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$.
Proof.
We first prove the returned solution has a cost at most $(1+\epsilon)Z^{*}$. Let $\pi^{1}$ denote the schedule for $I^{\prime}$ when we consider the rejection type and the assignment of large jobs associated with the schedule $\pi^{\prime}$ stated in Lemma 15. By the linear program LP(11), the cost of $\pi^{1}$ is bounded by the cost of $\pi^{\prime}$ if we ignore the jobs in $T_{2}\cup S_{3}$. Note that each job $J_{j}\in S_{3}$ is safe and thus by Definition 7, $e_{j}\leq\Delta$. Since we reject the jobs in $S_{3}$, by Lemma 20, the increased rejection penalty is at most $(m+1)\Delta$. Moreover, each job $J_{j}\in T_{2}$ is tiny and thus by Definition 16, $p_{j}\leq\Delta$. Therefore, the increment on the makespan caused by the jobs in $T_{2}$ is at most $2\Delta$. By Lemma 15 and Eqs.(9), (8), the cost of $\pi^{1}$ is at most
$(1+\frac{\epsilon}{2})Z^{*}(I)+(m+3)\Delta\leq(1+\frac{\epsilon}{2})Z^{*}(I)+\frac{\epsilon}{4}\leq(1+\epsilon)Z^{*}(I).$
When we substitute back the original rejection penalties, the cost does not increase. Therefore, the cost of the solution produced is at most $(1+\epsilon)Z^{*}(I)$.
By Lemmas 14 and 19, the number of distinct rejection types and large job assignments are both fixed constants. The linear program is polynomial-time solvable in $m,n$. Therefore, Approx2 is an EPTAS for $Pm\mid rej,U,r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$. ∎
@@ 5 Conclusions
We study the scheduling with rejection problem where the total machine cost does not exceed a given upper bound $\hat{U}$. We assume the number $m$ of machines is a fixed constant and the problem is denoted as $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$. We present a simple $2$-approximation algorithm for $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$ and then further improve it to an EPTAS. Our EPTAS enumerates all possible rejection types and all possible large job assignment types to determine the rejection decisions for all risky jobs and the schedules of all accepted large risky jobs. Finally, we establish a linear program to deal with the safe and tiny jobs.
One possible future research direction is to design an FPTAS for $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$. Note that when $m$ is part of the input, the problem, denoted as $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$, becomes strongly NP-hard. Therefore, another research direction is to extend our results to present a PTAS for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}+\sum_{J_{j}\in R}e_{j}$. Moreover, we currently assume the energy cost rates are uniform and it is natural to think about the non-uniform energy cost rates.

### v2

@@ 4.3 Determine the rejection decisions for safe jobs
$|S_{3}|\leq\frac{1}{\gamma^{2}}+2$.
@@
Since the number of constraints of LP(11) is $|S|+\frac{1}{\gamma^{2}}+2$, there are at most $|S|+\frac{1}{\gamma^{2}}+2$ positive values in the basic solution. By Eq.(12), it is easy to see the number of positive values is at least $|S_{1}|+|S_{2}|+2|S_{3}|=|S|+|S_{3}|$. Therefore, we have $|S|+|S_{3}|\leq|S|+\frac{1}{\gamma^{2}}+2$ and thus the lemma is proved. ∎
We reject all jobs in $S_{2}\cup S_{3}$ and accept all jobs in $S_{1}$ and in other words, we have determined the rejection decisions for all jobs. Let $A^{\prime}$ be the set of the accepted jobs in $\pi^{\prime}$ and $A$ be the set of the accepted jobs in our solution. One clearly sees that $A^{\prime}\cap B=A\cap B$ by Lemma 6 and $y_{j}=1$ only for each $J_{j}\in A^{\prime}\cap S$ is a feasible solution for LP(11). Therefore, we have
$\sum_{J_{j}\in A\cap S}p_{j}\leq\sum_{J_{j}\in A^{\prime}\cap S}p_{j}\mbox{ and }|A\cap SL_{i}|\leq\ell_{i},\mbox{ for every $i=1,\ldots,1/\gamma^{2}$}.$ (13)
It follows that our solution is feasible, that is, the total processing time of the accepted jobs is no more than $\hat{U}$. The following lemma shows that there exists a schedule for the accepted jobs such that the cost is bounded by $(1+\frac{\epsilon}{6})Z^{*}(I)+\frac{3}{2}\gamma+\frac{3}{2}\gamma^{3}$.
@@
When the rejection decisions are determined by our previous steps, there exists a schedule for the accepted jobs such that the total cost is at most $(1+\frac{\epsilon}{6})Z^{*}(I)+\frac{3}{2}\gamma+\frac{3}{2}\gamma^{3}$.
@@
Recall that the cost comes from two parts: the makespan of the accepted jobs and the total rejection penalty of the rejected jobs.
We construct a schedule $\pi^{\prime\prime}$ for the set of accepted jobs $A$ as follows:
  • 1.
    Recall that $A^{\prime}\cap B=A\cap B$ and thus we schedule each job in $A\cap B$ on the same machine as in $\pi^{\prime}$.
  • 2.
    Given a machine, we greedily schedule the jobs in $A\cap SL_{i}$ on the machine until the number of assigned jobs is exactly the same as that of $\pi^{\prime}$ or we run out of jobs.
  • 3.
    Finally, we greedily assign the jobs in $A\cap S\cap T$ to the least load machine.
We discuss the makespan of $\pi^{\prime\prime}$. For ease of presentation, we let $C_{\max}$ be the makespan of $\pi^{\prime}$ and clearly, $C_{\max}\geq\frac{1}{m}\sum_{J_{j}\in A^{\prime}}p_{j}$. If the makespan of $\pi^{\prime\prime}$ is determined by a job in $S\cap T$, then since the processing time of each job in $T$ is at most $\gamma$, its makespan is at most $\frac{1}{m}\sum_{J_{j}\in A}p_{j}+\gamma$. By Eq.(13), $A^{\prime}\cap B=A\cap B$ and $C_{\max}\geq\frac{1}{m}\sum_{J_{j}\in A^{\prime}}p_{j}$, the makespan of $\pi^{\prime\prime}$ is at most $C_{\max}+\gamma$. Otherwise, the makespan of $\pi^{\prime\prime}$ is determined by a job in $B$ or some $SL_{i}$. For every two jobs $J_{j},J_{k}\in SL_{i}$, by Definition 6, $p_{j}-p_{k}\leq\gamma^{2}$. By the second item of Lemma 6, there are at most $\frac{1}{\gamma}$ jobs in $S\cap L=\cup_{i}SL_{i}$ assigned to a same machine. Therefore, the increment in the makespan is bounded by $\frac{1}{\gamma}\times\gamma^{2}=\gamma$. In conclusion, the makespan of $\pi^{\prime\prime}$ is at most $C_{\max}+\gamma$.
We next consider the rejection penalty of the rejected jobs. The rejection penalty of $\pi^{\prime}$ is at least $\frac{K}{2}\gamma^{3}$ by the definition of $K$. Moreover, by Lemma 7, the rejection penalty of our solution is at most $\frac{K+1}{2}\gamma^{3}+(\frac{1}{\gamma^{2}}+2)\times\frac{1}{2}\gamma^{3}$. The increment in the rejection cost is at most $\frac{1}{2}\gamma+\frac{3}{2}\gamma^{3}$. Finally, by Lemma 6, we know that the cost of $\pi^{\prime\prime}$ is at most $(1+\frac{\epsilon}{6})Z^{*}(I)+\frac{3}{2}\gamma+\frac{3}{2}\gamma^{3}$. The proof is completed. ∎
Note that when the accepted job set $A$ is fixed, the problem is reduced to the classic multiprocessor scheduling problem $P\mid\mid C_{\max}$. Therefore, we can call a PTAS of $P\mid\mid C_{\max}$ for the job set $A$ and get the final solution. We present the high-level description of the whole algorithm in Algorithm 2.
Algorithm 2 Algorithm Approx2 for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$
1:Input: An instance $J=\{J_{1},\ldots,J_{n}\}$, $m$ identical machines and a fixed constant $0<\epsilon<1$;
2:Call Approx1 in Algorithm 1 on the instance II to obtain cost $Z(I)\leq 2Z^{*}(I)$;
3:use $Z(I)$ to normalize instance II so that $\frac{1}{2}\leq Z^{*}(I)\leq Z(I)=1$ as in Eq.(8);
4:set up $\gamma=\frac{\epsilon}{24}$ as in Eq.(9) and partition jobs into risky and safe/large and tiny as in Definition 4;
5:set up $\delta=\frac{\epsilon\gamma^{3}}{12}$ as in Eq.(10) and construct the restricted instance $I^{\prime}$as in Definition 5;
6:partition the risky jobs in $I^{\prime}$ into at most $\frac{1}{\delta}$ subsets $B_{i}$ and the jobs that are both safe and large into at most $\frac{1}{\gamma^{2}}$ subsets $SL_{i}$ as in Definition 6;
7:for each rejection type $(\ell_{1},\ldots,\ell_{1/\gamma^{2}},k_{1},\ldots,k_{1/\delta},K)$ as in Definition 7 do
8:  reject the first $k_{i}$ jobs in $B_{i}$ for every $i=1,\ldots,1/\delta$ and accept the other risky jobs;
9:  construct the linear program as in LP(11) and compute a basic solution;
10:  partition the safe jobs into three sets $S_{1},S_{2},S_{3}$ as in Eq.(12);
11:  accept the jobs in $S_{1}$ and reject the jobs in $S_{2}\cup S_{3}$;
12:  call a PTAS of $P\mid\mid C_{\max}$ on the accepted jobs;
13:end for
14:return the feasible solution with the minimum cost for $I^{\prime}$;
15:replace back the original rejection penalty of each job in the selected solution.
Theorem 2.
Algorithm Approx2 is a PTAS for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$.
@@
We first prove the returned solution has a cost at most $(1+\epsilon)Z^{*}(I)$. Let $\pi^{1}$ denote the schedule for $I^{\prime}$ when we consider the rejection type associating with the schedule $\pi^{\prime}$ stated in Lemma 6. Note that the PTAS used in Line 12 will cause another $1+\frac{\epsilon}{3}$ loss on the makespan. By Lemma 6 and Eqs.(9), (8), the cost of $\pi^{1}$ is at most
$(1+\frac{\epsilon}{3})\left((1+\frac{\epsilon}{6})Z^{*}(I)+\frac{3}{2}\gamma+\frac{3}{2}\gamma^{3}\right)$
$\leq$ $(1+\frac{\epsilon}{3})\left((1+\frac{\epsilon}{6})Z^{*}(I)+2\gamma\right)$ (By $3\gamma^{3}\leq\gamma$)
$=$ $(1+\frac{\epsilon}{3})\left((1+\frac{\epsilon}{6})Z^{*}(I)+\frac{\epsilon}{12}\right)$ (By Eq.(9))
$\leq$ $(1+\frac{\epsilon}{3})^{2}Z^{*}(I)$ (By Eq.(8))
$=$ $(1+\epsilon)Z^{*}(I).$ . (By $\epsilon<1$)
When we replace back the original rejection penalties, the cost is not increasing. Therefore, the cost of the outputted solution is at most $(1+\epsilon)Z^{*}(I)$.
By Lemma 5, the number of distinct rejection types is $N_{r}=(\frac{m}{\gamma}+1)^{1/\gamma^{2}}\times(\frac{2}{\gamma^{3}}+1)^{1/\delta}\times(\frac{2}{\gamma^{3}}+\frac{8}{\gamma^{2}}+1)$, which is polynomial in $m$. Therefore, the total running time is polynomial in $m,n$ since solving LP(11) and the PTAS for $P\mid\mid C_{\max}$ can be done in polynomial time in $m,n$. In conclusion, Approx2 is a PTAS for $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. ∎
@@ 5 Conclusions
We study the scheduling with rejection problem where the total machine cost does not exceed a given upper bound $\hat{U}$. We assume the number $m$ of machines is a part of the input and the problem is denoted as $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$. We present a simple $2$-approximation algorithm for $P\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ and then further improve it to a PTAS. Our PTAS enumerates all possible rejection types to determine the rejection decisions for all risky jobs and finally, we establish a linear program to deal with the rejection decisions of safe jobs.
One possible future research is to design an FPTAS for $Pm\mid rej,\hat{U},r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ where $m$ is a fixed constant. Moreover, we currently assume the energy cost rates are uniform and it is natural to think about the non-uniform energy cost rates.